The width of a rectangle is 3 inches less than the length. The perimeter is 46 inches. Find the length and the width.
The length is
inches and the width is
inches.

Answer by Cromlix(4381)

Hi there,
Make the length = 'x'
width = x - 3
Perimeter of a rectangle
= 2(length) + 2(width)
46 inches = 2(x) + 2(x - 3)
46 inches = 2x + 2x - 6
collect like terms:
4x = 46 + 6
4x = 52
x = 13 inches
Length is 13 inches
Width is 10 inches
Hope this helps :-)